OpenStudy (anonymous):

i'm not sure what to do here Kevin had 3/4 yard of string. he used 1/3 of the string for a school project. How many yards of string did he use for his project?

OpenStudy (anonymous):

(1/3) * (3/4) = 3/12 = 1/4 so its 1/4

OpenStudy (anonymous):

so multiply or subtract??

OpenStudy (anonymous):

multiply, im positive

OpenStudy (anonymous):

its 1/4

OpenStudy (anonymous):

y multiply?

OpenStudy (anonymous):

its either 1/3 but if thats not in the choices its 1/4

OpenStudy (anonymous):

it is they both are

OpenStudy (anonymous):

O.O oh no...wait a sec

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

XD im trying to make sure you get it correct...this seems to be a subtracting problem...so its most likely, if im correct, 1/3

OpenStudy (anonymous):

yep, it says he "used" so i think you should go with 1/3

OpenStudy (anonymous):

lol thank you

OpenStudy (anonymous):

your welcome, please tell me if I was correct

OpenStudy (anonymous):

ok i will :)

OpenStudy (anonymous):

I'll be here waiting :)

OpenStudy (anonymous):

oh ok but there are still more questions in my test i'll send u a message if u want?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

if you need help...just tag me :D

OpenStudy (anonymous):

ok thanks so much for the help :)

OpenStudy (anonymous):

your welcome